OpenStudy (anonymous):

Solve. 2x2 + x = 15.

OpenStudy (msmr):

I'm assuming that means \[2x ^{2} + x = 15\]?

OpenStudy (anonymous):

yes

OpenStudy (hoblos):

2x2 + x = 15 2x2 + x -15 =0 (2x-5)(x+3)=0 x= 5/2 OR x=-3

OpenStudy (msmr):

okay, you'll need to use the quadratic formula: \[ \frac{-b \pm \sqrt{b ^{2}} - 4ac}{2a}\]

OpenStudy (anonymous):

\[2x^2+x-15=0\] \[(2x-5)(x+3)=0\] x=2,5 or -3

OpenStudy (hoblos):

@sfozturk61 2&5 are not solutions 5/2 & -3 are the only solutions **

OpenStudy (anonymous):

Rewrite as 2x^2 +6x - 5x -15. Now factor by grouping 2x(x +3) - 5(x + 3). Factors to (2x- 5)(x+3)= 0 x=5/2 and -3

OpenStudy (anonymous):

@hoblos I mean 2,5 so 2 and one half = 5/2